OpenStudy (anonymous):

what is the equivalent fraction for 3/4 up to 10

OpenStudy (sandra):

ok I think you're asking for the first 10 equivalent fractions for 3/4

OpenStudy (sandra):

in that case, remember you can multiply any number by "1" and get the same number - including fractions. it may just look different

OpenStudy (sandra):

so 1/1, 2/2, 3/3, 4/4 are all different ways of writing 1

OpenStudy (sandra):

so to get the first 10 fractions , you want to multiply 3/4 by 2/2, 3/3, 4/4, etc... So the first equivalent fraction is: 3/4 x 2/2 = 2x3/4/2 = 6/8

OpenStudy (sandra):

you can get the rest!